OpenStudy (anonymous):

Find the slope of the line containing the given pair of points. If the line is undefined, state so. Please be sure to show all of your work. (1, -9) and (-2, -3)

OpenStudy (callisto):

slope = (y2-y1)/(x2-x1) = [(-3)-(-9)] / [(-2)-1] = 6/ (-3) = -2